# DeepMath-103K: A Large-Scale, Challenging, Decontaminated, and Verifiable Mathematical Dataset for Advancing Reasoning

### v1

@@ Abstract
The capacity for complex mathematical reasoning is a key benchmark for artificial intelligence. While reinforcement learning (RL) applied to LLMs shows promise, progress is significantly hindered by the lack of large-scale training data that is sufficiently challenging, possesses verifiable answer formats suitable for RL, and is free from contamination with evaluation benchmarks. To address these limitations, we introduce DeepMath-103K, a new, large-scale dataset comprising approximately 103K mathematical problems, specifically designed to train advanced reasoning models via RL. DeepMath-103K is curated through a rigorous pipeline involving source analysis, stringent decontamination against numerous benchmarks, and filtering for high difficulty (primarily Levels 5-9), significantly exceeding existing open resources in challenge. Each problem includes a verifiable final answer, enabling rule-based RL, and three distinct R1-generated solutions suitable for diverse training paradigms like supervised fine-tuning or distillation. Spanning a wide range of mathematical topics, DeepMath-103K promotes the development of generalizable reasoning. We demonstrate that models trained on DeepMath-103K achieve significant improvements on challenging mathematical benchmarks, validating its effectiveness. We release DeepMath-103K publicly to facilitate community progress in building more capable AI reasoning systems.
(a) Difficulty Levels
(b) Improvement of SFT and RL using DeepMath-103K
Figure 1: We introduce DeepMath-103K, a large-scale dataset of challenging mathematical problems. DeepMath-103K is more challenging compared to existing datasets (Figure a). All problems have both verified final answer and multiple solution paths, supporting a broad spectrum of training paradigms that substantially improves reasoning performance (Figure b).
@@ 1 Introduction
Mathematical reasoning stands as a cornerstone capability for advanced artificial intelligence, serving as a critical proving ground for models aiming to emulate sophisticated human problem-solving (Kojima et al., 2022; Wei et al., 2022). Recent strides, particularly leveraging reinforcement learning (RL) with large language models (LLMs), have demonstrated significant potential in tackling complex mathematical problems demanding logical deduction, symbolic manipulation, and multi-step reasoning (Jaech et al., 2024; Guo et al., 2025; Team, 2024; xAI, 2025; Google, 2025). Notably, methods like RL-Zero (Guo et al., 2025), which employ online RL guided by binary rewards from verifiable answers, have surpassed traditional supervised fine-tuning approaches.
Despite this promise, the advancement of these powerful RL techniques is critically hampered by a bottleneck: the scarcity of suitable training data. Existing mathematical datasets (Hendrycks et al., 2021b; Cobbe et al., 2021a; Yu et al., 2024; Toshniwal et al., 2024) often fall short in several key aspects. They may lack the extreme difficulty needed to push the boundaries of current models, miss the verifiable answer format essential for rule-based RL reward schemes, suffer from contamination with standard evaluation benchmarks (compromising evaluation integrity), or are simply insufficient in scale, particularly concerning highly challenging problems. While human-annotated datasets tailored for RL (Wang et al., 2024; Face, 2025; Hu et al., 2025; Luo et al., 2025b; Yu et al., 2025; Dang & Ngo, 2025; Albalak et al., 2025) provide valuable insights, they often struggle with scale and capturing the extreme difficulty characteristic of competitive mathematics, which is necessary for training state-of-the-art reasoners.
To bridge this critical gap, we introduce DeepMath-103K, a novel, large-scale mathematical dataset specifically engineered to accelerate the development of advanced reasoning models via reinforcement learning. DeepMath-103K is distinguished by its high concentration of challenging mathematical problems, significantly surpassing the difficulty distribution prevalent in existing open datasets (Figure 1a). Our dataset construction methodology incorporates rigorous data decontamination against a comprehensive suite of benchmarks (Figure 4) to ensure trustworthy evaluation, alongside filtering for problems predominantly at high difficulty levels ($\geq 5$).
Crucially, every problem within DeepMath-103K features a verifiable final answer, directly enabling the application of rule-based reward functions in RL frameworks. Furthermore, each problem is accompanied by three distinct R1-generated solutions (Guo et al., 2025), offering rich data for diverse training paradigms such as supervised fine-tuning, reward modeling, or model distillation (Figure 2). The dataset covers a broad spectrum of mathematical topics (Figure 3), ranging from foundational concepts to advanced subjects, thereby promoting the development of generalizable reasoning abilities. Comprising approximately 103K problems — including 95K curated challenging examples (Levels 5-10) and 8K supplementary problems (Levels 3-5) — DeepMath-103K represents a significant contribution toward equipping the community with the resources required to train highly proficient mathematical reasoners.
We validate the effectiveness of DeepMath-103K by demonstrating that models trained on it achieve substantial performance gains on several demanding mathematical reasoning benchmarks. This work furnishes a vital, openly accessible resource, tackling the urgent need for large-scale, challenging, verifiable, and clean mathematical data essential for propelling the next generation of AI reasoning systems.
Our main contributions are as follows:
  • We design and release DeepMath-103K, a novel large-scale mathematical dataset specifically curated for training advanced reasoning models via reinforcement learning, characterized by its high difficulty, verifiable answers, multiple diverse solutions per problem, and rigorous decontamination.
@@
  • We detail a meticulous data curation pipeline, encompassing source analysis, extensive decontamination against standard benchmarks for evaluation integrity, difficulty filtering, and robust answer verification.
  • We demonstrate the effectiveness of DeepMath-103K by showing that models trained on it achieve significant improvements on multiple challenging mathematical reasoning benchmarks, particularly using RL-Zero techniques enabled by the dataset’s structure.
@@ 2 Overview of DeepMath-103K
Figure 2: An example data sample from the DeepMath-103K dataset, illustrating its components.
@@
  • Final Answer: A verifiable final answer, crucial for enabling rule-based reward functions in RL settings.
  • Difficulty: A numerical difficulty score, which facilitates techniques like difficulty-aware training (e.g., curriculum learning) or adaptive compute allocation based on problem complexity (Wang et al., 2025; Chen et al., 2024).
  • Topic: A hierarchical topic classification for the problem, enabling topic-specific analysis or training.
  • R1 Solutions: Three distinct reasoning paths generated by the DeepSeek-R1 model (Guo et al., 2025), suitable for diverse training paradigms.
@@
DeepMath-103K includes mathematical problems spanning difficulty levels 3 through 10. The dataset comprises 95K challenging problems (levels 5-10) specifically curated for this research, augmented with an additional 8K problems (levels 3-5) sourced from SimpleRL (Zeng et al., 2025b) to ensure broader difficulty coverage. For comparison, we analyzed and labeled the difficulty levels of several existing datasets commonly used for RL training in this domain: Open-RS (Dang & Ngo, 2025), DAPO-17K (Yu et al., 2025), DeepScaleR-Preview (Luo et al., 2025b), ORZ-129K (Hu et al., 2025), and Open-R1 (Face, 2025). Figure 1a illustrates the difficulty distributions across these datasets. As depicted, DeepMath-103K exhibits a significantly more challenging problem distribution, containing a substantially higher proportion of problems at difficulty levels 5 and above compared to the other benchmark datasets. This focus on higher difficulty is intended to push the reasoning limits of current models.
@@
A key characteristic of DeepMath-103K, alongside its high difficulty, is its extensive topical diversity across the mathematical landscape. We systematically categorized each problem using a hierarchical topic structure, following the methodology established by Gao et al. (2024). As illustrated in Figure 3, this classification reveals that DeepMath-103K draws problems from a multitude of core mathematical areas. The dataset’s scope ranges from fundamental topics such as Prealgebra and Plane Geometry to sophisticated domains like Abstract Algebra (including Group Theory and Field Theory) and advanced Calculus (covering Differential Equations and Applications of Integrals, among others). This broad topical foundation ensures that models trained on DeepMath-103K are exposed to a rich variety of mathematical concepts and problem-solving paradigms, thereby fostering the development of more robust and widely generalizable reasoning skills.
Rigorous Data Decontamination
DeepMath-103K was constructed exclusively using the training splits of existing open-source datasets, with careful avoidance of any known test set materials. However, our preliminary analysis revealed that these source datasets exhibit concerning levels of contamination with problems from commonly used evaluation benchmarks. As illustrated in Figure 4, the contamination rates (defined as the percentage of benchmark test samples found within the raw data pool aggregated from source training splits) were notably high: reaching 90% for AIME24 and AMC23, 76.6% for MATH500, 35.7% for Minerva Math, and 33.6% for OlympiadBench. Recognizing that these benchmarks are frequently employed for model evaluation, DeepMath-103K underwent a rigorous decontamination procedure. This process systematically identified and removed problems that overlap with these standard evaluation sets, ensuring the integrity of future benchmark results obtained using models trained on DeepMath-103K.
Suitable for Diverse Training Paradigms
A core advantage of DeepMath-103K is its comprehensive data structure with both a verified final answer and multiple solution paths, supporting a broad spectrum of training paradigms and research methodologies.
  • Supervised Fine-Tuning: By providing three distinct R1-generated solutions for each problem, DeepMath-103K enables the creation of rich supervised training corpora. These solutions offer multiple valid approaches to the same question, allowing a model to learn diverse problem-solving strategies. In contrast to datasets that present only a single correct path, the variety in DeepMath-103K’s solutions helps LLMs better generalize to unseen problems.
  • Model Distillation: Advanced teacher–student paradigms often rely on having multiple labeled solution paths for individual questions. Through DeepMath-103K’s three solution trajectories, a larger teacher model can effectively impart varied problem-solving styles to a smaller student model, enhancing the student’s coverage of different reasoning heuristics and strategies.
  • Rule-based Reinforcement Learning (e.g., RL-Zero): The availability of a verifiable final answer in each problem allows for straightforward reward assignment, as models can be directly evaluated on whether their predicted result matches the correct solution. This binary feedback, essential for RL-based methods such as RL-Zero (Guo et al., 2025), fosters deeper reasoning by encouraging improvements specifically targeted at correctness.
  • Reward Modeling: With multiple valid solution paths, it becomes possible to design reward models that differentiate between high- and low-quality reasoning steps or compare alternative solution routes. This not only refines policy gradients in RL frameworks but also aids in re-ranking or scoring candidate solutions in a multi-step decoding pipeline (xAI, 2025).
Taken together, these features make DeepMath-103K exceptionally flexible for cutting-edge research on AI-driven mathematical reasoning, supporting both direct RL with correctness-based rewards and more nuanced learning frameworks that benefit from multiple, validated solution paths.
@@ 3 Construction of DeepMath-103K
Figure 6: Difficulty distributions of various open mathematical reasoning datasets considered as potential sources.
@@
To identify data sources rich in challenging problems, we began by analyzing the landscape of existing open mathematical reasoning datasets. These datasets utilize diverse collection methodologies. For instance, datasets such as MetaMathQA (Yu et al., 2024), dart-math-hard (Tong et al., 2024), and OpenMathInstruct-2 (Toshniwal et al., 2024) primarily focus on augmenting problems and solutions derived from established datasets like GSM8K (Cobbe et al., 2021a) and MATH (Hendrycks et al., 2021b). In contrast, datasets like NuminaMath-CoT (LI et al., 2024), MMIQC (Liu et al., 2024), and WebInstructSub (Yue et al., 2024) source content more broadly from the web, gathering materials such as exercises and discussions from online platforms (e.g., Math Stack Exchange).
We follow Gao et al. (2024) to estimate the difficulty distribution of these potential source datasets, as shown in Figure 6. This analysis revealed distinct patterns: datasets derived mainly from GSM8K and MATH (MetaMathQA, dart-math-hard, OpenMathInstruct-2), along with NuminaMath-CoT, exhibited distributions heavily skewed towards lower difficulty levels (levels 1-5). Conversely, datasets sourced more broadly from web content, specifically MMIQC and WebInstructSub, displayed significantly flatter distributions with a larger proportion of problems in the mid-to-high difficulty range (levels 5-9). Based on this finding, we selected MMIQC and WebInstructSub as our primary data sources due to their higher concentration of challenging problems. We also included NuminaMath-CoT to enhance the topical diversity of the initial collection. After applying basic filtering, this selection process yielded a raw pool of 2,869K questions.
@@
MATH500 Sorry to bother you today, but I came across this question: There are finitely many primes $p$ for which the congruence $8x\equiv 1\pmod{p}$ has no solutions $x$. Determine the sum of all such $p$. I first thought the answer was $0,$ since they didn’t say that $x$ had to be an integer, but apparently, it did. I don’t know how to proceed from here, any solutions? Thanks for taking the time to read this! There are finitely many primes $p$ for which the congruence $8x\equiv 1\pmod{p}$ has no solutions $x$. Determine the sum of all such $p$.
Table 1: Examples of contamination detected between the raw data pool and benchmarks using semantic comparison. Colors highlight conceptual or textual similarities leading to flagging.
Stage 2: Data Decontamination.
As indicated by the high contamination rates observed in common benchmarks (Figure 4), a rigorous data decontamination process was crucial for ensuring the integrity of DeepMath-103K. We performed decontamination against a comprehensive suite of mathematics and STEM benchmarks, including MATH (Hendrycks et al., 2021b), AIME (MAA, a), AMC (MAA, b), Minerva Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), Omni-MATH (Gao et al., 2024), MathOdyssey (Fang et al., 2024), GAOKAO (Zhong et al., 2023), JEEBench (Arora et al., 2023), MMLU-STEM (Hendrycks et al., 2021a), CMATH (Wei et al., 2023), OlympicArena (Huang et al., 2024), GSM8K (Cobbe et al., 2021a), and GPQA (Rein et al., 2024). We adopted the decontamination method proposed by Toshniwal et al. (2024) to effectively remove potential paraphrases of benchmark questions. This involved the following steps:
  1. 1.
  2. 2.
    Each candidate question was then compared against its top-$k$ retrieved benchmark examples using an LLM-Judge (Llama-3.3-70B-Instruct (Grattafiori et al., 2024)) to determine if they constituted identical questions or paraphrases. To mitigate positional bias, we evaluated each pair twice, swapping the order of the questions, resulting in $2k$ comparisons per candidate. If any of these $2k$ comparisons indicated a potential paraphrase or duplicate, the candidate question was discarded.
This semantic approach aims to identify not only exact duplicates but also near-duplicates and paraphrased questions that might otherwise overlap with evaluation sets.
Table 1 illustrates the effectiveness of this semantic decontamination compared to simple lexical matching. For instance, the AIME24 and AMC23 examples show that questions conveying similar mathematical concepts were correctly flagged as contaminated, despite differences in numerical values and phrasing. The MATH500 example demonstrates the detection of a near-identical question embedded within conversational text. These cases underscore the capability of our method to identify subtle conceptual overlaps, thereby yielding a dataset largely free from leakage relative to common evaluation benchmarks.
@@
6 A square has one side lying on the line y=2x17y=2x-17 and two other vertices on the parabola $y=x^{2}$. Determine the minimum possible area of the square.”
7 Determine the sequence $s(k,n)$, which represents the number of sides of the intersection of a unit-radius regular polygon $P_{k}$ with $k$ sides and a rotating unit-radius regular polygon $P_{n}$ with $n\geq k$ sides, as the angle of rotation $\theta$ varies from 0 to $2\pi$. Provide the sequence $s(k,n)$ for all $n\geq k$.
@@
Stage 3: Difficulty Filtering.
Prior work Zeng et al. (2025a) highlights the importance of aligning RL training data difficulty with the target model’s reasoning capabilities, noting that powerful models benefit significantly from exposure to highly challenging problems. Building on this insight, our curation process for DeepMath-103K focuses on selecting problems that represent a significant reasoning challenge. To quantify difficulty, we adopted the approach detailed in Omni-MATH (Gao et al., 2024). We assigned a difficulty level to each decontaminated problem by prompting GPT-4o based on the annotation guidelines provided by the Art of Problem Solving (AoPS) [^1] https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings. To ensure a robust estimate, we queried GPT-4o six times for each problem and averaged the resulting ratings to determine its final difficulty level. Subsequently, we applied a strict filtering criterion, retaining only those problems with an estimated difficulty level of 5 or higher. Table 2 showcases examples of geometry problems that passed this filtering stage, illustrating how increasing difficulty levels often correlate with greater conceptual depth and reasoning complexity.
@@
The availability of verifiable final answers is crucial for enabling rule-based reward mechanisms in RL, which helps mitigate reward hacking and has been instrumental in training successful reasoning models like DeepSeek-R1 (Guo et al., 2025). However, reliably constructing such answers presents two primary challenges:
@@
    Some question types, such as mathematical proofs or open-ended explorations, inherently lack a unique, easily verifiable final result.
@@
    Question Formatting and Filtering: We utilized GPT-4o to process the raw questions. Problem types inherently unsuitable for verification (e.g., proofs) were discarded. Questions phrased conversationally were automatically rewritten into a standardized format seeking a single, specific numerical or symbolic answer.
@@
    Answer Verification via Consistency Check: For questions successfully passing the above step, we generated three distinct solution paths using the DeepSeek-R1 model. A rule-based verifier then extracted the final answer from each of these generated solutions, as well as from the original source solution (when available). We enforced strict consistency: only problems where all extracted final answers were identical were retained in the final dataset.

### v2

@@ Abstract
Reinforcement learning (RL) with large language models shows promise in complex reasoning. However, its progress is hindered by the lack of large-scale training data that is sufficiently challenging, contamination-free and verifiable. To this end, we introduce DeepMath-103K, a large-scale mathematical dataset designed with high difficulty (primarily levels 5-9), rigorous decontamination against numerous benchmarks, and verifiable answers for rule-based RL reward. It further includes three distinct R1 solutions adaptable for diverse training paradigms such as supervised fine-tuning (SFT). Spanning a wide range of mathematical topics, DeepMath-103K fosters the development of generalizable and advancing reasoning. Notably, models trained on DeepMath-103K achieve state-of-the-art results on challenging mathematical benchmarks and demonstrate generalization beyond math such as biology, physics and chemistry, underscoring its broad efficacy.
(a) Difficulty Levels of different datasets.
(b) Pass@1 Accuracies on AIME25.
Figure 1: (a) DeepMath-103K is challenging compared to existing datasets. (b) Results of DeepMath series models under zero RL and RL setting using DeepMath-103K.
@@ 1 Introduction
Reinforcement learning (RL) with large language models (LLMs) has demonstrated significant potential in complex mathematical reasoning (Guo et al., 2025; Hu et al., 2025; Zeng et al., 2025a; Liu et al., 2025). Despite this promise, the effective advancement of RL is constrained by existing training data. While numerous datasets are available, they fall short in several key aspects crucial for training advanced reasoning models: (1) insufficient difficulty (Figure 1a) to push the boundaries of current models (Dang & Ngo, 2025; Yu et al., 2025; Luo et al., 2025; Face, 2025; Hu et al., 2025), (2) contamination with standard benchmarks (appendix A), (3) a lack of verifiable answers essential for RL with verifiable rewards (RLVR) (Guo et al., 2025; Cobbe et al., 2021; Hendrycks et al., 2021b; Yu et al., 2024), or (4) an inadequate combination of these critical aspects at scale. Furthermore, many of existing datasets represent the recombination and filtration of common sources (such as AIME (MAA, a)) which contain already well-formatted data, thus lacking a substantial influx of novel and diverse problems from more varied but less structured sources (Dang & Ngo, 2025; Yu et al., 2025; Luo et al., 2025; Face, 2025; Hu et al., 2025).
To bridge this gap, we introduce DeepMath-103K, a large-scale mathematical dataset tailored for advancing reasoning via RLVR. DeepMath-103K distinguishes itself through several key features.
  • Challenging Problems: DeepMath-103K features a high concentration of challenging mathematical problems, with a difficulty distribution skewed towards higher levels ($\geq 5$) compared to existing open resources (Figure 1a).
  • Rigorous Decontamination: To ensure trustworthy evaluation, DeepMath-103K underwent a rigorous decontamination process against a comprehensive suite of benchmarks.
  • Verifiable Answers and Diverse Solutions: To enable rule-based reward functions in RLVR, every problem in DeepMath-103K includes a verifiable final answer that has been validated for easy extraction and verification via rules. Each problem is further enriched with three distinct R1 solutions (Guo et al., 2025), supporting diverse training paradigms such as SFT.
Beyond these core features, DeepMath-103K also differentiates itself in its raw data acquisition. The prevalent trend in existing open datasets often recombines readily available and well-formatted problems from common sources such as AIME (MAA, a). This approach does not create new problems, but re-collect existing ones, which leads to significant overlaps among different datasets. Recognizing the potential limitations and eventual exhaustion of common resources, DeepMath-103K draws its content from more diverse but less structured sources, notably including discussions from Math StackExchange [^1] https://math.stackexchange.com. The raw content from these sources is informal discourse and lacking a standard format. After a rigorous curation pipeline that transformed these discussions into a well-structured QA format, DeepMath-103K is characterized by its unique problem variety and diversity compared to existing datasets.
Consequently, models trained on DeepMath-103K achieve state-of-the-art (SOTA) results (Figure 1b):
  • Zero RL Training: Starting from the Qwen-2.5-(Math)-7B (Team, 2024), DeepMath-Zero-(Math)-7B shows pass@1 improvements of +12.7 (+23.0) on AIME24 and +12.1 (+19.1) on AIME25, establishing new SOTA performance.
  • RL Training: Initialized from instruction-tuned models, DeepMath variants also show substantial gains. DeepMath-1.5B, starting from R1-Distill-Qwen-1.5B (Guo et al., 2025), achieves pass@1 accuracy improvements of +7.9 on AIME24 and +6.0 AIME25. DeepMath-Omn-1.5B, built upon OpenMath-Nemotron-1.5B (Moshkov et al., 2025), reaches new SOTA pass@1 accuracies of 64.0 on AIME24 and 57.3 on AIME25, surpassing o1-mini (63.6 on AIME24) and low effort o3‑mini (60.0 on AIME24).
  • Generalizable Reasoning beyond Math: DeepMath series models also generalizes their reasoning abilities to broader domains, achieving best GPQA-Diamond (Rein et al., 2024) scores on biology, physics, and chemistry compared to the baselines.
These results underscore the value of DeepMath-103K as a resource for developing advanced reasoning models with broad applicability. The remainder of this paper is organized as follows:
  • § 2 presents an overview of DeepMath-103K, including its format, difficulty distribution, and topic covered;
@@
  • § 3 details the data curation pipeline to construct DeepMath-103K, encompassing source analysis, decontamination, difficulty filtering, and robust answer verification
  • § 4 trains, evaluates and analyzes DeepMath series models that trained on DeepMath-103K.
To foster future research, we have released the DeepMath-103K dataset, along with the code and model weights, hoping to enable further exploration of advanced reasoning techniques and the development of robust and generalizable machine intelligence.
@@ 2 Overview of DeepMath-103K
Figure 2: A data sample from DeepMath-103K.
@@
  • Final Answer: A verifiable final answer, crucial for rule-based reward functions in RLVR.
  • Difficulty: A numerical difficulty score, which facilitates techniques like difficulty-aware training (e.g., curriculum learning) or adaptive compute allocation based on problem complexity (Wang et al., 2025b; Chen et al., 2024).
  • Topic: A hierarchical topic classification for the problem, enabling topic-specific analysis.
  • R1 Solutions: Three distinct reasoning paths generated by the DeepSeek-R1 model (Guo et al., 2025), suitable for diverse training paradigms such as SFT.
@@
DeepMath-103K includes mathematical problems spanning difficulty levels 3 through 9. The core of the dataset consists of 95K challenging problems (levels 5-9) specifically curated for this research. To ensure broader difficulty coverage, this is augmented with an additional 8K problems (levels 3-5) sourced from SimpleRL (Zeng et al., 2025b). For comparison, we analyzed and labeled the difficulty levels of several existing datasets commonly used for RLVR training in math domain: Open-RS (Dang & Ngo, 2025), DAPO-17K (Yu et al., 2025), DSR-Preview (Luo et al., 2025), SITLL-3-RL (Chen et al., 2025), ORZ-129K (Hu et al., 2025), and Open-R1 (Face, 2025). Figure 1a illustrates the difficulty distributions across these datasets. As depicted, DeepMath-103K exhibits a significantly more challenging problem distribution, containing a substantially higher proportion of problems at difficulty levels 5 and above compared to the other benchmark datasets. This focus on higher difficulty is intended to push the reasoning limits of current models.
Rigorous Data Decontamination
DeepMath-103K was constructed exclusively using the training splits of existing open resources, with careful avoidance of any known test set materials. However, our preliminary analysis revealed that these source data exhibits alarmingly high levels of contamination with problems from commonly used evaluation benchmarks. As illustrated in Figure 3, the contamination rates (defined as the percentage of benchmark test samples found within our raw data pool) are notably high: reaching 90% for AIME24 and AMC23, 76.6% for MATH500, 35.7% for Minerva Math, and 33.6% for OlympiadBench. Recognizing that these benchmarks are frequently employed for model evaluation, DeepMath-103K underwent a rigorous decontamination procedure. This process systematically identified and removed problems that overlap with these standard evaluation sets, ensuring the integrity and reliability of future benchmark results obtained using models trained on DeepMath-103K.
Broad Topical Diversity
@@
Complementing its high difficulty and data integrity, a key characteristic of DeepMath-103K is its extensive topical diversity spanning the mathematical landscape. We categorized each problem using a hierarchical topic structure, following the methodology from Gao et al. (2024). As illustrated in Figure 4, this classification reveals that DeepMath-103K draws problems from a multitude of core mathematical areas. Its scope ranges from fundamental topics such as Prealgebra and Plane Geometry to sophisticated domains like Abstract Algebra (including Group Theory and Field Theory) and advanced Calculus (covering Differential Equations and Applications of Integrals, among others). This broad and deep topical foundation ensures that models trained on DeepMath-103K are exposed to a rich variety of mathematical concepts and problem-solving paradigms, thereby fostering the development of more robust and widely generalizable reasoning skills.
Data Novelty and Uniqueness
As mentioned in § 1, DeepMath-103K sources mostly from math forum, rather than common resources frequently adopted by other datasets. To evaluate the data novelty and uniqueness of DeepMath-103K, we performed the following analysis for all the datasets:
  1. 1.
    We first embedded all the samples using paraphrase-multilingual-MiniLM-L12-v2.
  2. 2.
    Samples with an embedding similarity greater than 0.98 were considered as the same samples.
Viewing each dataset as a set of embeddings, Figure 5 presents the number of unique elements in each set and the corresponding set sizes. DeepMath-103K contains 82.81K problems that are not found in others. This stark contrast highlights the data novelty and uniqueness of DeepMath-103K. We also plot their embedding distribution after t-SNE in Figure 6. ORZ-129K (Hu et al., 2025), Open-R1 (Face, 2025), SITLL-3-RL (Chen et al., 2025), DSR-Preview (Luo et al., 2025), and DAPO-17K (Yu et al., 2025), though curated independently, show very similar embedding distribution, while DeepMath-103K exhibits a distinctly different pattern. This observation supports our claim that existing datasets overlap with each other because of using common data sources and further demonstrate the data novelty and uniqueness of DeepMath-103K.
Figure 5: Unique and non-unique problem counts in DeepMath-103K compared to other datasets.
(a) DeepMath-103K
(b) ORZ-129K
(c) Open-R1
(d) STILL-3-RL
(e) DSR-Preview
(f) DAPO-17K
Figure 6: Embedding distributions of different datasets after t-SNE.
@@ 3 Construction of DeepMath-103K
Figure 8: Difficulty distributions of various open mathematical datasets considered as potential sources.
@@
To identify data sources rich in challenging problems, we began by analyzing the landscape of existing open mathematical reasoning datasets designed for SFT. These datasets utilize diverse collection methods. For instance, datasets such as MetaMathQA (Yu et al., 2024), dart-math-hard (Tong et al., 2024), and OpenMathInstruct-2 (Toshniwal et al., 2024a) primarily focus on augmenting problems and solutions derived from established datasets like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). In contrast, datasets like NuminaMath-CoT (LI et al., 2024), MMIQC (Liu et al., 2024), and WebInstructSub (Yue et al., 2024) source content more broadly from the web, gathering materials such as exercises and discussions from online platforms (e.g., Math Stack Exchange). We follow Gao et al. (2024) to estimate the difficulty distributions of these potential source datasets, as shown in Figure 8, which reveals distinct patterns: datasets derived from GSM8K and MATH (MetaMathQA, dart-math-hard, OpenMathInstruct-2), along with NuminaMath-CoT, exhibited distributions heavily skewed towards lower difficulty levels (levels 1-5). Conversely, datasets sourced more broadly from web content, specifically MMIQC and WebInstructSub, displayed significantly flatter distributions with a larger proportion of problems in the mid-to-high difficulty range (levels 5-9). Based on this finding, we selected Math StackExchange subsets from MMIQC and WebInstructSub as our primary data sources due to their higher concentration of challenging problems. We also included NuminaMath-CoT to enhance the topical diversity of the initial collection. After applying basic filtering, this selection process yielded a raw pool of 2,869K questions.
Stage 2: Data Decontamination.
As indicated by the high contamination rates observed in common benchmarks (Figure 3), a rigorous data decontamination process was crucial for ensuring the integrity of DeepMath-103K. We performed decontamination against a comprehensive suite of mathematics and STEM benchmarks, including MATH (Hendrycks et al., 2021b), AIME (MAA, a), AMC (MAA, b), Minerva Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), Omni-MATH (Gao et al., 2024), MathOdyssey (Fang et al., 2024), GAOKAO (Zhong et al., 2023), JEEBench (Arora et al., 2023), MMLU-STEM (Hendrycks et al., 2021a), CMATH (Wei et al., 2023), OlympicArena (Huang et al., 2024), GSM8K (Cobbe et al., 2021), and GPQA (Rein et al., 2024). We adopted the decontamination method proposed by Toshniwal et al. (2024a):
  1. 1.
  2. 2.
    Each candidate question was then compared against its top-$k$ retrieved benchmark examples using an LLM-Judge (Llama-3.3-70B-Instruct (Grattafiori et al., 2024)) to determine if they constituted identical questions or paraphrases. If any of these comparisons indicated a potential paraphrase or duplicate, the candidate question was discarded.
Table 1 illustrates the effectiveness of semantic decontamination compared to simple lexical matching. This approach aims to identify not only exact duplicates but also near-duplicates and paraphrased questions that might otherwise overlap with evaluation sets.
Table 1: Examples of contamination detected between the raw data pool and benchmarks using semantic comparison. Colors highlight conceptual or textual similarities.
@@
Stage 3: Difficulty Filtering.
@@
6 A square has one side lying on the line $y=2x-17$ and two other vertices on the parabola $y=x^{2}$. Determine the minimum possible area of the square.”
7 Determine the sequence $s(k,n)$, which represents the number of sides of the intersection of a unit-radius regular polygon $P_{k}$ with $k$ sides and a rotating unit-radius regular polygon $P_{n}$ with $n\geq k$ sides, as the angle of rotation $\theta$ varies from 00 to $2\pi$. Provide the sequence $s(k,n)$ for all $n\geq k$.
@@
Zeng et al. (2025a) highlights the importance of aligning RL training data difficulty with the target model’s reasoning capabilities, noting that powerful models benefit significantly from exposure to challenging problems. Building on this insight, our curation process for DeepMath-103K focuses on selecting problems that represent a significant reasoning challenge. To quantify difficulty, we adopted the approach detailed in Omni-MATH (Gao et al., 2024). We assigned a difficulty level to each decontaminated problem by prompting GPT-4o based on the annotation guidelines provided by the AoPS [^2] https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings. To ensure a robust estimate, we queried GPT-4o six times for each problem and averaged the resulting ratings to determine its final difficulty level. Subsequently, we applied a strict filtering criterion, retaining only those problems with an estimated difficulty level of 5 or higher. Table 2 showcases examples of geometry problems that passed this filtering stage, illustrating how increasing difficulty levels often correlate with greater conceptual depth and reasoning complexity.
@@
The availability of verifiable final answers is crucial for enabling rule-based reward in RLVR, which helps mitigate reward hacking and has been instrumental in training successful reasoning models like DeepSeek-R1 (Guo et al., 2025). However, reliably constructing such answers presents two primary challenges:
@@
    Some open-ended questions inherently lack a easily verifiable final answer.
@@
    Question Filtering and Formatting: We used GPT-4o to process the raw questions. Problem types inherently unsuitable for verification were discarded. Questions phrased conversationally were rewritten into a standardized format seeking a single, specific numerical or symbolic answer.
@@
    Answer Verification via Consistency Check: For questions successfully passing the above step, we generated three distinct solution paths using DeepSeek-R1. A rule-based verifier then extracted the final answer from each of these generated solutions, as well as from the original source solution (when available). We enforced strict consistency: only problems where all extracted final answers were identical were retained in the final dataset.